OpenStudy (anonymous):

Can someone tell me the equation to find the volume of a circle? I forgot-_-

OpenStudy (amistre64):

a circle is 2 dimensional and does not pertain to volume, which is 3 dimensional

OpenStudy (amistre64):

the "area" of a circle is: pi r^2

OpenStudy (anonymous):

my bad its a container... read the question wrong

OpenStudy (amistre64):

if the container is a cylindar, then its base area, times height

OpenStudy (amistre64):

if its a sphere ... a ball .... i believe we get to 4/3 pi r^3

OpenStudy (anonymous):

ok think you could just help me out with the problem?

OpenStudy (amistre64):

type up the problem and lets have a look at it

OpenStudy (anonymous):

container 1 Radius 2 inches Volume 50.24 cubic inches container 2 Radius 3 inches Volume ? So i have to find the volume of container 2 heres my answer choices 113.04 169.56 254.34 75.36

OpenStudy (anonymous):

Based off of container 1 my guess is D but want to be sure.

OpenStudy (amistre64):

well, 32pi/3 is equal to about 33.51 so this doesnt seem to relate to a sphere are there pictures to go by? similar shapes perhaps?

OpenStudy (anonymous):

no thats why im so lost with this problem

OpenStudy (anonymous):

it does say this though in the beginning "A company makes similar cylindrical containers" so i guess theyre cylindrical containers

OpenStudy (amistre64):

or if its just a r to V ratio ... 2:50.24 relates to 1:25.12 3(1):3(25.12) 3:75.36

OpenStudy (anonymous):

its not a ratio but that is probably the easiest way to solve it thanks for showing me that!

OpenStudy (anonymous):

it still gives you the right answer so i think we could just use that... i just have to find the answer dont have to do it a certain way and show my work.

OpenStudy (amistre64):

let assume its a cylindar then, and that radius refers to a circular base: V = pi r^2 * h 50.24 = 4pi h h = 50.24/4pi V = 9pi * 50.24/4pi = 9 * 50.24/4 = 113.04

OpenStudy (amistre64):

so depending on the context of the problem, there are a few ways to try to approach it. but as is, its not very well defined to me

OpenStudy (anonymous):

yeah i agree. Thanks for the help anyways!

OpenStudy (anonymous):

can u helpme with my question now? http://openstudy.com/study#/updates/51c1c859e4b0bb0a4360320a